OpenStudy (anonymous):

how do i find the maximum and minimum points of a function? f(x) 3x^2+6X-10

OpenStudy (anonymous):

do i need to use a quadratic formula?

OpenStudy (callisto):

re-arrange the quadratic equation into the form f(x) = a(x-h)^2 +k If a>0, the curve opens upwards => a minimum point, and vice versa (h,k) is the coordinates of the vertex In your case, f(x) =3x^2+6X-10 = 3(x^2 +6x) -10 =3(x^2 + 6x +3^2 - 3^2) -10 = 3 (x^2 +6x +9) - 3x3x3 -10 = 3(x+3)^2 - 37 So, can you work out the answer?

OpenStudy (mimi_x3):

You can differentiate; it's easier I suppose.

OpenStudy (anonymous):

it was already differentiated from f(x)=x3+3x^2-10 x

OpenStudy (anonymous):

a book i was using said to use a quadratic formula.i remember using a calculator to solve it once,but i cant remember.

OpenStudy (callisto):

Then you should out f'(x) =0 and solve the equation After that, draw a table and see the change of slope If its change is like -ve -> +ve , then it is a minimum point, and vice versa

OpenStudy (mimi_x3):

Then set it to zero. \[3x^2+6x-10=0\] You can use the quadratic formula if you want.

OpenStudy (callisto):

Sorry i didn't know that you have differentiated it.

OpenStudy (mimi_x3):

You can also find the second derivative to determine the nature.